OpenStudy (anonymous):

4(x – b) = x (1 point)

OpenStudy (anonymous):

4(x-b) = x 4x-4b =x -4b=-3x b=3/4x

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

No Problem